Assume Cylinder A and Cone B are the same height and the bases have the same radius. If A has a volume of 18π cm3, what is the volume of B? (round to nearest whole number)
1
Expert's answer
2020-05-06T16:48:23-0400

Let "V_A" be the volume of cylinder, "V_B" be the volume of cone. We know that "V_A = \pi r_A^2h_A", "V_B = \pi r_B^2h_B/3" , where "h" is the height and "r" is the radius. "h_A = h_B" and "r_A = r_B" by conditions of the problem. We have

"V_B = \pi r_b^2h_b/3 = \pi r_A^2h_A/3 = V_A/3= \\=18\pi/3=6\pi= 18.84cm^3 \approx 19 cm^3."

Answer: "19cm^3".